OpenStudy (b.r):

Math? Fan & medal? 4 questions?

OpenStudy (b.r):

1. t + 27 = 21 t = –6 t = 47 t = 6 t = –47 2. 24 = m – 32 m = –8 m = 56 m = 768 m = –56 3. Tamara has eaten 1,400 calories for breakfast and lunch. She has 800 more calories, c, that she can eat for dinner. Write and solve an equation to find Tamara’s daily caloric intake. 1,400 – 800 = c; c = 600 1,400 + 800 = c; c = 2,200 1,400 – c = 800; c = 2,200 1,400 + 800 = c; c = 600 4. Steven’s old cell phone cost $125.00 less than his new phone. His old cell phone cost $155.00. Write and solve an equation to find the cost of Steven’s new cell phone. n – 125 = 155; n = $30 155 – 125 = n; n = $30 125 + n = 155; n = $280 n – 125 = 155; n = $280

OpenStudy (anonymous):

Starting with number one, subtract 27 from both sides

OpenStudy (anonymous):

For 1 , subtract the like terms For 2 , add the like terms